OpenStudy (anonymous):

solve using the addition principle. y-(-3/4)=-5/6 please show work

OpenStudy (anonymous):

y+3/4 = -5/6 y = -5/6-3/4 = -(5/6+3/4) = -19/12

OpenStudy (vannah):

Y=-19/20 YOU ARE GOING TO CROSS MULTIPLY-3/4 -5/6 FIVE TIMES FOUR =20 SIX TIMES 3 IS 18 AND THE Y IS AN INVISBILE ONE THREE TIMES SIX IS NINETEEN PLUS 1 =-19/20

OpenStudy (anonymous):

vannah, that is wrong. the correct answer is what I posted

jhonyy9 (jhonyy9):

vannah you need to check the common denominator for 6 and 4 what is 12 and not is 20